elevator weighting 600kg is pulled upward by cable with an acceleration of 4.0m/s what is the tension in the cable?what is the tension when the elevator is accelerating downward at 4.0m/s,but is stilling moving upward

Grade:12th pass

1 Answers

Saksham Rastogi
6 years ago
Let us write an equation for the elevator: 
 
T – mg = ma 
=> T – 600 X 10 = 600 X 4
=> T = 2400 + 6000
=> T = 8400 (Ans.)